# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### 20 Jan 2024 v1

@@ 3.3 Dataset Statistics
Finally, the statistical distributions presented in Figures 23678 and 9 demonstrate that our dataset maintains very similar distributions across all the predefined dataset splits. First, this consistency highlights the reliability of our entire dataset and its predefined cross-validation folds and test set, showing that all the categories in all the proposed tasks keep the same frequency distribution in all splits. Second, this observation validates our annotation process as it proves the efficiency of our designed annotation protocols and our multiple stages of data validation and correction.
@@
We include further statistical analysis in Appendix B of the Supplementary Material. This section provides information about the total number of instances in each task of the GraSP dataset, the transition probabilities for both phase and steps, the intra-task categories co-occurrences for the instruments and actions tasks, and the detailed inter-task categories co-occurrences for instruments-phases, instrument-steps, and instruments-actions in GraSP. Our previously presented statistics and the additional statistics in the Supplementary Material prove the intrinsic complementarity among the proposed tasks in GraSP. We previously established the relation between phases as steps where their per-class frequency and duration distributions demonstrated similar patterns related to the semantic meaning of the categories. Regarding this, the most durable phase categories encompass many of the most durable step categories, and the least durable phases encompass many short steps. Likewise, there is a clear relationship between the surgical instrument and action categories, where many actions can be performed only by specific instruments. Thus, the frequency of different instruments affects the frequencies of actions. Additionally, the co-occurrence matrixes demonstrate an evident correspondence between instruments and phases or steps as, intuitively, different instruments are more present or completely absent in different phases and steps.
@@ 3.4 Evaluation Metrics
To evaluate the performance of phase and step recognition, we keep the originally proposed metric in Valderrama et al. (2022): the mean Average Precision ($mAP$) per frame. We chose this metric as it is a standard and robust metric commonly used for frame-wise evaluation of action recognition (Caba Heilbron et al. (2015)). We use the implementation of Python’s Scikit-learn library for the calculation of this metric 555sklearn.metrics.average_precision_score.
@@
Additionally, we maintain the usual semantic segmentation evaluation metrics we introduced in González et al. (2020). These metrics include the Mean Intersection over Union (mIoU), Intersection over Union (IoU), and Mean Class Intersection over Union (mcIoU). We kept these metrics as they are considerably stringent metrics that have been the standard evaluation method for the instrument segmentation task in multiple works. We utilize the implementation provided in Ayobi et al. (2023)  [^7] https://github.com/BCV-Uniandes/MATIS.
@@ 4 Transformers for Actions, Phases, Steps and Instrument Segmentation
We propose the Transformers for Actions, Phases, Steps, and Instrument Segmentation (TAPIS) model, a generalized architecture designed to tackle all the proposed tasks in the GraSP benchmark. Figure 10 portrays the general architecture of TAPIS. Our method utilizes a localized instrument segmentation baseline applied on independent keyframes that acts as a region proposal network and provides pixel-precise masks and their corresponding segment embeddings. Further, our model uses a global video feature extractor on time windows centered on a keyframe to compute a class embedding and a sequence of spatio-temporal embeddings. A frame classification head uses the class embedding to classify the middle frame of the time window into a phase or a step, and a region classification head interrelates the global spatio-temporal features with the localized region embeddings for atomic action prediction or instrument region classification. In this section, we explain the details of our proposed architecture.
@@ 4.1 Instrument Segmentation Baseline
We refer the reader to the original Mask2Former (Cheng et al. (2022)) and MATIS (Ayobi et al. (2023)) papers for further details on the Mask2Former architecture and its adaptation for surgical instrument segmentation. Furthermore, our TAPIS model utilizes Mask2Former’s region proposals two-fold. First, as in MATIS, to improve temporal consistency in instrument segmentation by reclassifying region proposals utilizing temporal context provided by a video feature extractor. Second, to handle the implicit need for instrument localization in atomic action detection. Thereby, TAPIS uses the corresponding segment embeddings of the proposed regions as inputs for a region classification head that enhances the segment classification process and predicts the atomic actions performed by the proposed instruments. We provide further insights about our region classification module in Section 4.3.
@@ 5 Experiments and Results
We present our overall test results in Table 4. Based on the experimental setup presented in Valderrama et al. (2022), we compare the performance of the best TAPIS model with a CNN-based baseline. For this alternative baseline, we employ SlowFast as the central video feature extractor and use ISINet (based on Mask R-CNN (He et al. (2017))) as the region proposal network for short-term tasks. We used the exact implementation details presented in Valderrama et al. (2022). We call this model SlowFast in Table 4. We also compare TAPIS with an alternative implementation using Video Swin Transformer (VST) (Liu et al. (2022)) as the video feature extractor and maintaining our proposed region classification head; we call this model TAPIS-VST in Table 4. The results in Table 4 demonstrate TAPIS’s and VST-TAPIS’s superiority over the CNN-based counterpart. This behavior proves that the inherent sequence-processing capability of transformers allows for a superior comprehension of temporal progressions, and their attention mechanisms result in better recognition of instruments, anatomical cues, and other relevant visual contexts in surgical videos, leading to higher performance in all the proposed semantic tasks. Likewise, TAPIS surpasses TAPIS-VST in most tasks and achieves comparable instrument segmentation performance with significantly fewer parameters. The outperformance of TAPIS over TAPIS-VST in phases, steps, and actions recognition proves that our global video feature extractor is optimal and sufficient to understand the temporal progressions of surgical procedures. Although TAPIS-VST yields a better instrument region classification performance due to its finer-grained spatial processing and its increased amount of parameters, it still surpasses TAPIS by less than a 0.5% relative improvement with more than twice the parameters, hence highlighting TAPIS’s satisfactory capacity.
@@
Moreover, Table 4 also compares TAPIS’ performance with the TAPIR model (Valderrama et al. (2022)). To ensure a fair comparison, we train TAPIR with its original implementation details, parameters, and box proposal architecture (Deformable DETR) on the training set of GraSP and test its performance on the test set. This comparison reveals that TAPIS significantly outperforms TAPIR across all tasks in the GraSP benchmark, especially in the novel and highly challenging atomic action detection task, where TAPIS achieves more than a 53% relative improvement over TAPIR’s results. The performance difference proves the potential of our instrument segmentation annotations for short-term tasks and underscores the impact of our technical improvements and methodological experimentation. Notably, TAPIR exhibits a slight decrease in performance for short-term tasks compared to the originally reported metrics in PSI-AVA [^11] https://github.com/BCV-Uniandes/TAPIR, which we attribute to the refining and curation of our dataset that eliminated biases and included annotations for challenging tiny instrument segments that were previously unannotated. The superiority of TAPIS over TAPIR and all alternative baselines positions it as state-of-the-art in all tasks of our GraSP dataset.
@@ 5.1.1 Instruments and Actions Presence Recognition
Given that GraSP includes instrument localization annotations, we expand our model evaluation to cover simpler and coarser tasks, such as recognizing the presence of instruments and associated actions. For this purpose, we adapt the dataset by transforming annotations into single binary vectors of dimensions $[1,N]$ for multi-label classification in each frame, where $N$ represents the number of classes (7 for instruments presence and 14 for actions presence). To enable TAPIS to execute this task, we modified the model output to consolidate predictions from each predicted region into a single classification vector. Since it is our best region proposal method, we employ our instrument segmentation baseline with a SwinL backbone as the region proposal model.
We compare TAPIS’s performance with two alternative baselines to validate our data and model. First, an alternative implementation of TAPIS without region proposal features that predicts instruments and actions presence directly from the video feature extractor. Second, with Rendezvous (Nwoye et al. (2022)), a state-of-the-art model for recognizing the presence of actions and instruments. This model predicts triplets in the form of $<$Instrument, Verb, Target$>$, where verb categories indicate the actions of each instrument in the keyframe, treated as analogous to the action categories in GraSP. To adapt Rendezvous for instrument presence recognition, we utilize all instrument categories, along with the “null verb” and “null target” categories, forming the triplets. Similarly, for the action presence task, we employ all action categories, along with the “instrument” and “null target” categories, to create the triplets. Additionally, we modify Rendezvous’ outputs according to the multi-label nature of GraSP actions. This adjustment included generating triplets for all actions present within each keyframe. Importantly, we evaluated the models without pretraining on surgical datasets to ensure a fair comparison while preserving the pretraining in COCO of our region proposal network. The evaluation metric used for assessing performance was the $mAP$.
@@ 5.1.2 Interactions Recognition in Endovis 2018
We employ DINO (Zhang et al. (2023)) as our instrument detector and use the proposed $mAP@0.5IoU_{box}$ metric for evaluation. For a fair comparison, we evaluate the predictions made by the other models for each instrument (excluding the kidney category) as a conventional detection problem.
We compare our method with MSLRGR-GISFSG (Seenivasan et al. (2022)), which simultaneously performs semantic instrument segmentation and interaction detection utilizing a graph reasoning approach. To compare accurately with this method, which relies on ground-truth bounding boxes for inference, we assume perfect bounding box locations for instruments, as this model does not perform direct detection. For interactions, we used the model’s predicted class for each box. We assign the bounding box class for instruments based on the instrument mask predicted within it.
@@ 5.1.3 Gesture Recognition in RARP45
We further validate our methodology using the RARP-45 public gesture recognition benchmark. Given the absence of a public test set in the original RARP-45 dataset, we create our splits, using 27 videos for training and 9 videos for evaluation. To ensure the integrity of the splits, we maintained a consistent class distribution with the original training set. Specifically, for the test split, we select videos 2, 3, 4, 6, 7, 9, 14, 17, 23, 28, 32, and 33, while we use the remaining videos for the training split. Finally, we extracted the frames of RARP45’s original videos using a frame rate of six frames per second.
Following our previous experimental framework, we compared our method with its alternative versions using SlowFast and VST as video feature extractors. All experiments employed a window of 16 frames with a stride of 1. As shown in Table 7, our TAPIS model outperforms the alternative baselines, with all transformer-based methods surpassing the CNN model. This behavior demonstrates our model’s robustness by achieving outstanding performance in tasks analogous to the ones proposed in GraSP yet differently formulated. These results are consistent with the findings in the GraSP dataset, thus reinforcing the effectiveness of our approach in handling video-level tasks and proving the reliability of our results.
@@ 5.1.4 Phase and Step Recognition in MISAW
To further validate our results in phases and steps recognition, we tested our models on the MISAW dataset (Huaulmé et al. (2021)), which is analogous in long-term tasks to GraSP. We again follow the experimental setup we defined for GraSP and compare the performance of the different explored architectures for GraSP (SlowFast and TAPIS-VST) in the phases and steps recognition benchmarks of the MISAW dataset.
@@ 5.2.1 Long-Term Tasks Experiments
In contrast, steps represent finer-grained time segments within these phases. Consequently, discerning these discrete steps requires a more detailed temporal context around the central frame to extract relevant features for step differentiation effectively. As detailed in Table 10, our analysis highlights that employing temporal windows sampled with shorter temporal stride (higher frame rates) results in higher performance in steps recognition. Increasing the temporal stride reduces performance, and excessively augmenting it yields the lowest $mAP$ values observed. We find that the optimal temporal window for step recognition is to keep a 16-frame window size and use a low 1-second temporal stride so the model captures a temporal window of 16 seconds centered on the frame of interest. Our optimal window design is consistently related to average step duration statistics from Figure 5, where step duration ranges mostly from 10 to 200 seconds approximately. Notably, steps exhibit a wider range of duration compared to phases. This variability implies that employing a uniform temporal window to identify all step categories poses a challenge, as certain instances might require a larger temporal context for accurate identification compared to others.
Additionally, we present more specific results of phases and steps recognition per-class performance in the Supplementary Material. Initially, the assessment of the model’s performance across diverse phase categories reveals differing levels of success in identifying individual surgical phases. Notably, phases 1-6 and 10 consistently attain relatively higher $mAP$ scores across all three methods, while categories 7 and 8 exhibit relatively lower scores. When correlating these observations with the dataset statistics (Figure 2), it is notable that the phases with lower $AP$ scores across the methods display a low representation in the dataset, as it is natural when studying imbalanced datasets. Additionally, the lower mean duration of these phases (Figure 4) might contribute to the model finding it more challenging to discern and accurately predict these phases, such as classes 7 and 8. This relation is related to the broader temporal stride that may not be optimal for shorter phase instances that span fewer consecutive frames.
@@
Model Backbone FLOPs (G) Params (M) $mAP@0.5IoU_{box}$
Faster R-CNN R50 180 42 80.88 ±$\pm 1.16$
Deformable DETR R50 173 40 87.33 ±$\pm 2.91$
DINO SwinL 310 218 90.93 ±$\pm 1,15$
Similarly, the model’s performance across various step categories reveals that steps 12 to 18 and 20, characterized by lower mean durations, exhibit relatively lower Average Precision (AP) scores across the evaluated methods. This lower performance might also stem from their shorter durations, making it more challenging for the model to discern and accurately predict these particular steps. Additionally, less represented step classes in the dataset, like 6, 7, and 11, tend to manifest lower $AP$ scores, and steps 0, 1, and 19, with higher representation, demonstrate comparatively better performance. However, the model’s performance was still low for some highly represented classes like 2,3 and 5. This effect may be due to high variability in the duration of these steps that we showcase in 5. Hence, the varying performance between classes is also related to the fixed temporal window design, which may not be optimally suited to capture the high variability in the duration of different step instances.

### 26 Jan 2024 v2

@@ 3.3 Dataset Statistics
The statistical distributions presented in Figures 23678 and 9 demonstrate that our dataset maintains similar distributions across all the predefined dataset splits. First, this consistency highlights the reliability of our entire dataset and its predefined cross-validation folds and test set, showing that all the categories in all the proposed tasks keep comparable frequency distributions in all splits. Second, this observation validates our annotation process as it proves the efficiency of our designed annotation protocols and our multiple stages of data validation and correction.
@@
We include further statistical analysis in C of the Supplementary Material. This section provides information about the total number of instances in each task of the GraSP dataset, the transition probabilities for both phase and steps, the intra-task categories co-occurrences for the instruments and actions tasks, and the detailed inter-task categories co-occurrences for instruments-phases, instrument-steps, and instruments-actions in GraSP. Our previously presented statistics and the additional statistics in the Supplementary Material prove the intrinsic complementarity among the proposed tasks in GraSP. We previously established the relation between phases as steps where their per-class frequency and duration distributions demonstrated similar patterns related to the semantic meaning of the categories. Regarding this, the most durable phase categories encompass many of the most durable step categories, and the least durable phases encompass many short steps. Likewise, there is a clear relation between the surgical instrument and action categories, where many actions can be performed only by specific instruments. Thus, the frequency of different instruments affects the frequencies of actions. Additionally, the co-occurrence matrixes demonstrate an evident correspondence between instruments and phases or steps as, intuitively, different instruments are more present or completely absent in different phases and steps.
@@ 3.4 Evaluation Metrics
To evaluate the performance of phase and step recognition, we keep the originally proposed metric in Valderrama et al. (2022): the mean Average Precision ($mAP$) per frame. We chose this metric as it is a standard and robust metric commonly used for frame-wise evaluation of action recognition (Caba Heilbron et al. (2015)). We use the implementation of Python’s Scikit-learn library 555sklearn.metrics.average_precision_score for the calculation of this metric.
@@
Additionally, we maintain the usual semantic segmentation evaluation metrics we introduced in González et al. (2020). These metrics include the Mean Intersection over Union (mIoU), Intersection over Union (IoU), and Mean Class Intersection over Union (mcIoU). We kept these metrics as they are highly stringent metrics that have been the standard evaluation methodology for the instrument segmentation task in multiple works. We utilize the implementation provided in Ayobi et al. (2023)  [^7] https://github.com/BCV-Uniandes/MATIS.
@@ 4 Transformers for Actions, Phases, Steps and Instrument Segmentation
We propose the Transformers for Actions, Phases, Steps, and Instrument Segmentation (TAPIS) model, a generalized architecture designed to tackle all the proposed tasks in the GraSP benchmark. Figure 10 portrays the general architecture of TAPIS. Our method utilizes a localized instrument segmentation baseline applied on independent keyframes that acts as a region proposal network and provides pixel-precise instrument masks and their corresponding segment embeddings. Further, our model uses a global video feature extractor on time windows centered on a keyframe to compute a class embedding and a sequence of spatio-temporal embeddings. A frame classification head uses the class embedding to classify the middle frame of the time window into a phase or a step, and a region classification head interrelates the global spatio-temporal features with the localized region embeddings for atomic action prediction or instrument region classification. In the following subsections, we explain the details of our proposed architecture.
@@ 4.1 Instrument Segmentation Baseline
We refer the reader to the original Mask2Former (Cheng et al. (2022)) and MATIS (Ayobi et al. (2023)) papers for further details on the Mask2Former architecture and its adaptation for surgical instrument segmentation. Furthermore, our TAPIS model utilizes Mask2Former’s region proposals two-fold. First, as in MATIS, to improve temporal consistency in instrument segmentation by reclassifying region proposals utilizing temporal context provided by a video feature extractor. Second, to handle the implicit need for instrument localization in atomic action detection. Thereby, TAPIS uses the corresponding segment embeddings of the proposed regions as inputs for a region classification head that enhances the segment classification process or predicts the atomic actions performed by the proposed instruments. We provide further insights about our region classification module in Section 4.3.
@@ 5 Experiments and Results
We present our overall test results in Table 4. Based on the experimental setup presented in Valderrama et al. (2022), we compare the performance of the best TAPIS model with a CNN-based baseline. For this alternative baseline, we employ SlowFast (Feichtenhofer et al. (2019)) as the central video feature extractor and use ISINet (González et al. (2020)) (based on Mask R-CNN (He et al. (2017))) as the region proposal network for short-term tasks. We use the exact implementation details of SlowFast presented in Valderrama et al. (2022). We call this model SlowFast in Table 4. We also compare TAPIS with an alternative implementation using Video Swin Transformer (VST) (Liu et al. (2022)) as the video feature extractor and maintaining our proposed region classification head; we call this model TAPIS-VST in Table 4. The results in Table 4 demonstrate TAPIS’s and VST-TAPIS’s superiority over the CNN-based counterpart. This behavior proves that the inherent sequence-processing capability of transformers allows for a superior comprehension of temporal progressions, and their attention mechanisms result in better recognition of instruments, anatomical cues, and other relevant visual contexts in surgical videos, leading to higher performance in all the proposed semantic tasks. Likewise, TAPIS surpasses TAPIS-VST in most tasks and achieves comparable instrument segmentation performance with significantly fewer parameters. The outperformance of TAPIS over TAPIS-VST in phases, steps, and actions recognition proves that our global video feature extractor is optimal and sufficient to understand the temporal progressions of surgical procedures. Although TAPIS-VST yields a better instrument region classification performance due to its finer-grained spatial processing and its increased amount of parameters, it still surpasses TAPIS by less than a 0.5% relative improvement with more than twice the parameters, hence highlighting TAPIS’s satisfactory performance.
@@
Moreover, Table 4 also compares TAPIS’ performance with the TAPIR model (Valderrama et al. (2022)). To ensure a fair comparison, we train TAPIR with its original implementation details, parameters, and box proposal architecture (Deformable DETR (Zhu et al. (2021))) on the training set of GraSP and test its performance on the test set. This comparison reveals that TAPIS significantly outperforms TAPIR across all tasks in the GraSP benchmark, especially in the novel and highly challenging atomic action detection task, where TAPIS achieves more than a 53% relative improvement over TAPIR’s results. The performance difference proves the potential of our instrument segmentation annotations for short-term tasks and underscores the impact of our technical improvements and methodological experimentation. Notably, TAPIR exhibits a slight decrease in performance for short-term tasks compared to the originally reported metrics in PSI-AVA [^11] https://github.com/BCV-Uniandes/TAPIR, which we attribute to the refining and curation of our dataset that eliminated biases and included annotations for challenging tiny instrument segments that were previously unannotated. The superiority of TAPIS over TAPIR and all alternative baselines positions it as state-of-the-art in all tasks of our GraSP dataset.
@@ 5.1.1 Instruments and Actions Presence Recognition
Given that GraSP includes instrument localization annotations, we expand our model evaluation to cover simpler and coarser tasks, such as recognizing the presence of instruments and associated actions. For this purpose, we adapt the dataset by transforming annotations into single binary vectors of dimensions $[1,N]$ for multi-label classification in each frame, where $N$ represents the number of classes (7 for instruments presence and 14 for actions presence). To enable TAPIS to execute this task, we modify the model output to consolidate predictions from each predicted region into a single classification vector. Since it is our best region proposal method, we employ our instrument segmentation baseline with a SwinL backbone as the region proposal model.
We compare TAPIS’s performance with two alternative baselines to validate our data and model. First, an alternative implementation of TAPIS without region proposal features that predicts instruments and actions presence directly from the video feature extractor. Second, with Rendezvous (Nwoye et al. (2022)), a state-of-the-art model for recognizing the presence of actions and instruments. This model predicts triplets in the form of $<$Instrument, Verb, Target$>$, where verb categories indicate the actions of each instrument in the keyframe, treated as analogous to the action categories in GraSP. To adapt Rendezvous for instrument presence recognition, we utilize all instrument categories, along with the “null verb” and “null target” categories, forming the triplets. Similarly, for the action presence task, we employ all action categories, along with the “instrument” and “null target” categories, to create the triplets. Additionally, we modify Rendezvous’ outputs according to the multi-label nature of GraSP actions. This adjustment included generating triplets for all actions present within each keyframe. Importantly, we train the models without pretraining on surgical datasets to ensure a fair comparison while preserving the pretraining in COCO of our region proposal network. The evaluation metric used for assessing performance was the $mAP$.
@@ 5.1.2 Interactions Recognition in Endovis 2018
We employ DINO (Zhang et al. (2023)) as our instrument detector and use the proposed $mAP@0.5IoU_{box}$ metric for evaluation. For a fair comparison, we evaluate the predictions made by the other models for each instrument (excluding the kidney category) as a conventional detection problem. We compare our method with MSLRGR-GISFSG (Seenivasan et al. (2022)), which simultaneously performs semantic instrument segmentation and interaction detection utilizing a graph reasoning approach. To compare accurately with this method, which relies on ground-truth bounding boxes for inference, we assume perfect bounding box locations for instruments, as this model does not perform direct detection. For interactions, we used the model’s predicted class for each box. We assign the bounding box class for instruments based on the instrument mask predicted within it.
@@ 5.1.3 Gesture Recognition in RARP45
We further validate our methodology using the RARP-45 public gesture recognition benchmark. Given the absence of a public test set in the original RARP-45 dataset, we create our splits, using 27 videos for training and 9 videos for evaluation. To ensure the integrity of the splits, we maintain a consistent class distribution with the original training set. Specifically, for the test split, we select videos 2, 3, 4, 6, 7, 9, 14, 17, 23, 28, 32, and 33, while we use the remaining videos for the training split.
Following our previous experimental framework, we compare our method with its alternative versions using SlowFast and VST as video feature extractors. All experiments employed a window of 16 frames with a stride of 1. As shown in Table 7, our TAPIS model outperforms the alternative baselines, with all transformer-based methods surpassing the CNN model. This behavior demonstrates our model’s robustness by achieving outstanding performance in tasks analogous to the ones proposed in GraSP yet differently formulated. These results are consistent with the findings in the GraSP dataset, thus reinforcing the effectiveness of our approach in handling video-level tasks and proving the reliability of our results.
@@ 5.1.4 Phase and Step Recognition in MISAW
To further validate our results in phases and steps recognition, we test our models on the MISAW dataset (Huaulmé et al. (2021)), which is analogous in long-term tasks to GraSP. We again follow the experimental setup we define for GraSP and compare the performance of the different explored architectures for GraSP (SlowFast and TAPIS-VST) in the phases and steps recognition benchmarks of the MISAW dataset.
@@ 5.2.1 Long-Term Tasks Experiments
In contrast, steps represent finer-grained time segments within these phases. Consequently, discerning these discrete steps requires a more detailed temporal context around the central frame to extract relevant features for step differentiation effectively. As detailed in Table 10, our analysis highlights that employing temporal windows sampled with shorter temporal stride (higher frame rates) results in higher performance in steps recognition. Increasing the temporal stride reduces performance, and excessively augmenting it yields the lowest $mAP$ values observed. We find that the optimal temporal window for step recognition is to keep a 16-frame window size and use a low 1-second temporal stride so the model captures a temporal window of 16 seconds centered on the frame of interest. Our optimal window design is consistently related to average step duration statistics from Figure 5, where step duration ranges mostly from 10 to 200 seconds approximately. Notably, steps exhibit a wider range of durations compared to phases. This variability implies that employing a uniform temporal window to identify all step categories poses a challenge, as certain instances might require a larger temporal context for accurate identification compared to others.
@@
Additionally, we present more specific results of phases and steps recognition per-class performance in the Supplementary Material. Initially, the assessment of the model’s performance across diverse phase categories reveals differing levels of success in identifying individual surgical phases. Notably, phases 1-6 and 10 consistently attain relatively higher $AP$ scores across all three methods, while categories 7 and 8 exhibit relatively lower scores. When correlating these observations with the dataset statistics (Figure 2), it is notable that the phases with lower $AP$ scores across the methods display a low representation in the dataset, as it is natural when studying long-tailed datasets. Additionally, the lower mean duration of these phases (Figure 4) might contribute to the model finding it more challenging to discern and accurately predict these phases, such as classes 7 and 8. This relation is related to the broader temporal stride that may not be optimal for shorter phase instances that span fewer consecutive frames.
@@
Model Backbone FLOPs (G) Params (M) $mAP@0.5IoU_{box}$
Faster R-CNN R50 180 42 80.88 ±$\pm 1.16$
Deformable DETR R50 173 40 87.33 ±$\pm 2.91$
DINO SwinL 310 218 90.93 ±$\pm 1,15$
@@
Similarly, the model’s performance across various step categories reveals that steps 12 to 18 and 20, characterized by lower mean durations, exhibit lower Average Precision ($AP$) scores across the evaluated methods. Additionally, less represented step classes in the dataset, like 6, 7, and 11, tend to manifest lower $AP$ scores, and steps 0, 1, and 19, with higher representation, demonstrate comparatively better performance. However, the model’s performance was still low for some highly represented classes like 2, 3 and 5. This effect may be due to the high variability in these steps’ duration, as shown in Figure 5. Once again, the varying performance between classes is related to the fixed temporal window design, which may not be optimally suited to capture the high inter and intra-class variability in the duration of different step instances.